# Power monoids and their arithmetic: a survey

### v1

@@ Abstract
To date, there is no satisfactory characterization of when $\mathcal{P}_{\mathrm{fin},1}(H)$ is UmF in the non-commutative case, and the situation is even more dire for HmFness. For partial progress, see Lemma 3.4(i), Theorem 3.6, and Proposition 4.5 in [13].
@@
The conjecture was first stated in [16, Sect. 5] for the special case $H=\mathbb{N}$, which, however, turns out to be equivalent to the general case. Indeed, any monoid $H$ containing an element of infinite order has a submonoid isomorphic to $\mathbb{N}$. By Proposition 3.2 and Theorem 4.1(ii), it follows that every length set of $\mathcal{P}_{\mathrm{fin},0}(\mathbb{N})$ is also a length set of $\mathcal{P}_{\mathrm{fin},0}(H)$.
@@
As for partial progress, it is known from [16, Propositions 4.8–4.10] that, for each $n\geq 2$, the singleton $\{n\}$, the interval $\llbracket 2,n\rrbracket$, and the set $\{2,n\}$ are length sets of $\mathcal{P}_{\mathrm{fin},0}(\mathbb{N})$. This remained the state of the art until recently, and some authors (including the present one) started to cast doubt on the conjecture. A breakthrough came in summer 2025, when Reinhart [34] proved that $\mathcal{P}_{\mathrm{fin},0}(\mathbb{N})$ is fully elastic: for every rational $q\geq 1$, there exists a length set $L$ of $\mathcal{P}_{\mathrm{fin},0}(\mathbb{N})$ such that $\max L=q\min L$.
We continue with a question reminiscent of the “Characterization Problem for Systems of Lengths” in the classical theory of factorizations [20, Section 6].
@@
Given a class $\mathcal{O}$ of finite monoids (for instance, the class of finite groups), determine whether there exists an integer $n\geq 1$ such that $H$ is isomorphic to $K$ for all $H,K\in\mathcal{O}$ with the property that (i) the systems of [minimal] length sets of $\mathcal{P}_{\mathrm{fin},1}(H)$ and $\mathcal{P}_{\mathrm{fin},1}(K)$ coincide and (ii) at least one of $H$ and $K$ has order greater than or equal to $n$.
The question is presented here for the first time. The answer is negative for the class of all finite monoids. Indeed, isomorphic monoids have the same systems of [minimal] length sets, and we gather from [47, Example 1.2(3)] that, for every $n\geq 1$, there exist non-isomorphic monoids $H$ and $K$ with $|H|=|K|=n$ whose reduced finitary power monoids are isomorphic. This naturally links Question 5.2 to Question 1.2.
We conjecture that the answer to Question 5.2 is positive for finite groups, for which Question 1.2 is known to have a positive answer [48]. Partial results for finite cyclic groups and minimal factorizations can be found in [5, Section 5], and there are at least a couple of non-trivial observations that are perhaps worth recording here.
First, we have from Proposition 3.2 and Theorem 4.1(ii) that the system of [minimal] length sets of a monoid $H$ is contained in the system of [minimal] length sets of $\mathcal{P}_{\mathrm{fin},1}(K)$ for every monoid $K$ into which $H$ embeds.
Second, [5, Proposition 4.12(i)] shows that, for every monoid $H$, the largest minimal length $\mu_{H}(X)$ of a set $X$ in $\mathcal{P}_{\mathrm{fin},1}(H)$ is bounded above by |X|1|X|-1. When $H$ is finite, it follows that the system of minimal length sets of $\mathcal{P}_{\mathrm{fin},1}(H)$ is contained in the power set of the interval $\llbracket 0,|H|-1\rrbracket$. Moreover, $H$ itself is an element of $\mathcal{P}_{\mathrm{fin},1}(H)$, and it seems reasonable to conjecture that $\mu_{H}(H)\to\infty$ as $|H|\to\infty$.
If confirmed, this latter conjecture would imply that, up to isomorphism, there are only finitely many pairs $(H,K)$ of finite monoids such that the systems of minimal lengths of $\mathcal{P}_{\mathrm{fin},1}(H)$ and $\mathcal{P}_{\mathrm{fin},1}(K)$ coincide.
@@
Let $\mathbb{Z}_{n}$ be the additive group of integers modulo $n$, where $n\geq 2$. In [12, Example 3.12(2)], Cossu and Tringali noted that
@@
On the other hand, we have already observed in the above that if $H$ contains an element of infinite order, then $\mathcal{P}_{\mathrm{fin},1}(H)$ has a divisor-closed submonoid isomorphic to $\mathcal{P}_{\mathrm{fin},0}(\mathbb{N})$. By [16, Propositions 4.8], this implies that
@@
One of the most intriguing aspects of power semigroups arises from the natural interplay between arithmetic and combinatorial properties. As a concrete example of this phenomenon (beyond the conjectures of Sárközy and Ostmann already mentioned in Section 1), let $H$ be a numerical monoid, and for all $k,l\in\mathbb{N}$ let $\alpha_{k,l}(H)$ denote the number of $k$-element atoms of $\mathcal{P}_{\mathrm{fin},0}(H)$ whose maximum is smaller than or equal to $l$. Experiments appear to support the following:
@@
The conjecture first appeared in an email from the author to Geroldinger and Gotti on January 20, 2023, where it was stated only for $H=\mathbb{N}$. It was later included as Problem 1 in the problem set of the “CrowdMath 2023: Arithmetic of Power Monoids” online collaborative project, organized by the same Gotti throughout 2023 and hosted on the Art of Problem Solving (AoPS) portal at
@@
As a first non-trivial step, Aggarwal et al. [1, Corollary 5.3] used a probabilistic approach to show that, in the special case $H=\mathbb{N}$, Conjecture 5.4 holds for almost all $l\in\mathbb{N}$ (here, “almost all” is understood in the sense of asymptotic density).

### v2

@@ Abstract
To date, there is no satisfactory characterization of when $\mathcal{P}_{\mathrm{fin},1}(H)$ is UmF in the non-commutative case, and the situation is even more dire for HmFness. For partial progress, see Lemma 3.4(i), Theorem 3.6, and Proposition 4.5 in [14]. By contrast, the situation is completely settled for atomic factorizations.
Theorem 4.19.
The following conditions hold:
  1. (i)
    $\mathcal{P}_{\mathrm{fin},1}(H)$ is UmF-atomic if and only if $H$ is trivial.
  2. (ii)
Proof.
These are Theorem 4.14 and Corollary 4.15 of [5], respectively. ∎
@@
The conjecture was first stated in [17, Sect. 5] for the special case $H=\mathbb{N}$, which, however, turns out to be equivalent to the general case. Indeed, any monoid $H$ containing an element of infinite order has a submonoid isomorphic to $\mathbb{N}$. By Proposition 3.2 and Theorem 4.1(ii), it follows that every length set of $\mathcal{P}_{\mathrm{fin},0}(\mathbb{N})$ is also a length set of $\mathcal{P}_{\mathrm{fin},0}(H)$.
@@
As for partial progress, it is known from [17, Propositions 4.8–4.10] that, for each $n\geq 2$, the singleton $\{n\}$, the interval $\llbracket 2,n\rrbracket$, and the set $\{2,n\}$ are length sets of $\mathcal{P}_{\mathrm{fin},0}(\mathbb{N})$. This remained the state of the art until recently, and some authors (including the present one) started to cast doubt on the conjecture. A breakthrough came in summer 2025, when Reinhart [36] proved that $\mathcal{P}_{\mathrm{fin},0}(\mathbb{N})$ is fully elastic: for every rational $q\geq 1$, there exists a length set $L$ of $\mathcal{P}_{\mathrm{fin},0}(\mathbb{N})$ such that $\max L=q\min L$.
That said, we continue with a question reminiscent of the “Characterization Problem for Systems of Lengths” in the classical theory of factorizations [21, Section 6].
@@
Given a class $\mathscr{C}$ of finite monoids (for instance, the class of finite groups), determine whether there exists an integer $n\geq 1$ such that if $H,K\in\mathscr{C}$ have the property that (i) the systems of minimal lengths of $\mathcal{P}_{\mathrm{fin},1}(H)$ and $\mathcal{P}_{\mathrm{fin},1}(K)$ coincide, and (ii) at least one of $H$ and $K$ has order greater than or equal to $n$, we have $H\cong K$.
The question is presented here for the first time. The answer is negative for the class of all finite monoids. Indeed, isomorphic monoids have the same systems of minimal lengths, and we gather from [48, Example 1.2(3)] that, for every $n\geq 1$, there exist non-isomorphic monoids $H$ and $K$ with $|H|=|K|=n$ whose reduced finitary power monoids are isomorphic. This naturally links Question 5.2 to Question 1.2.
We conjecture that the answer to Question 5.2 is positive for finite groups, for which it is known that Question 1.2 has a positive answer [50]. Partial results in the cyclic case appear in [5, Section 5], and there are a few more observations worth recording.
First, Conjecture 5.1 and Theorem 4.15 imply that any two non-trivial torsion-free monoids have the same system of lengths, consisting of $\{0\}$, $\{1\}$, and each non-empty finite subset of $\llbracket 2,\infty\rrbracket$ (cf. the comments under Conjecture 5.1). On the other hand, we have from [11, Example 2.2(1)] and [5, Proposition 4.7(v)] that every factorization in a unit-cancellative commutative monoid is minimal, and there are plenty of non-isomorphic torsion-free commutative monoids; for instance, it is folklore that two numerical monoids are isomorphic if and only if they are equal [26, Theorem 3]. All in all, this is part of the reason why Question 5.2 is restricted to finite monoids.
Second, it follows from Proposition 3.2 and Theorem 4.1(ii) that the system of minimal lengths of a monoid $H$ is contained in the system of minimal lengths of $\mathcal{P}_{\mathrm{fin},1}(K)$ for every monoid $K$ into which $H$ embeds.
Third, Theorem 4.17 shows that, for every monoid $H$, the largest minimal length $\mu_{H}(X)$ of a set $X$ in $\mathcal{P}_{\mathrm{fin},1}(H)$ is bounded above by |X|1|X|-1. When $H$ is finite, it follows that the system of minimal lengths of $\mathcal{P}_{\mathrm{fin},1}(H)$ lies in the power set of the interval $\llbracket 0,n-1\rrbracket$, where $n:=|H|$. Moreover, $H$ itself is an element of $\mathcal{P}_{\mathrm{fin},1}(H)$, and it is not difficult to establish that
$\mu_{H}(H)\geq\log_{2}(n).$ (5)
Indeed, assume $n\geq 2$ (otherwise the claim is obvious), and let $x_{1},\ldots,x_{n-1}$ be an enumeration of the non-identity elements of $H$. By Proposition 4.6, the word
$\mathfrak{a}:=\{1_{H},x_{1}\}\ast\cdots\ast\{1_{H},x_{n-1}\}$
is a factorization (into irreducibles) of $H$ in $\mathcal{P}_{\mathrm{fin},1}(H)$. On the other hand, if $\sigma$ is an injective map $\llbracket 1,\allowbreak k\rrbracket\to\allowbreak\llbracket 1,n-1\rrbracket$ for some $k\geq 1$, then for the word
$\{1_{H},\allowbreak x_{\sigma(1)}\}\ast\cdots\ast\{1_{H},x_{\sigma(k)}\}$
to be a factorization of $H$ it is necessary that $2^{k}\geq n$, i.e., $k\geq\log_{2}(n)$. Since every factorization of $H$ contains a subword that, up to a permutation of its letters, is a minimal factorization of $H$, this entails the desired inequality on $\mu_{H}(H)$.
In particular, Eq. (5) implies that, up to isomorphism, there are only finitely many pairs $(H,K)$ of finite monoids such that the systems of minimal lengths of $\mathcal{P}_{\mathrm{fin},1}(H)$ and $\mathcal{P}_{\mathrm{fin},1}(K)$ coincide (recall that, up to isomorphism, there exist finitely many monoids of order smaller than or equal to a fixed integer $N\geq 1$).
@@
As a starting point, let $\mathbb{Z}_{n}$ be the additive group of integers modulo $n$, where $n\geq 2$. In [13, Example 3.12(2)], Cossu and Tringali noted that
@@
On the other hand, we have already observed in the above that if $H$ contains an element of infinite order, then $\mathcal{P}_{\mathrm{fin},1}(H)$ has a divisor-closed submonoid isomorphic to $\mathcal{P}_{\mathrm{fin},0}(\mathbb{N})$. By [17, Propositions 4.8], this implies that
@@
To complete our list of problems, we now turn to one of the most intriguing features of power semigroups: the interplay between arithmetic and combinatorial properties.
As a concrete illustration of this phenomenon (beyond the conjectures of Sárközy and Ostmann already mentioned in Section 1), let $H$ be a numerical monoid, and for all $k,l\in\mathbb{N}$ let $\alpha_{k,l}(H)$ denote the number of $k$-element atoms of $\mathcal{P}_{\mathrm{fin},0}(H)$ whose maximum is less than or equal to $l$. Experiments appear to support the following:
@@
Prior to the present work, the conjecture had only been formulated in the special case $H=\mathbb{N}$. Originally proposed by the author in an email to Geroldinger and Gotti dated January 20, 2023, it was later included as Problem 1 in the problem set of the online collaborative project “CrowdMath 2023: Arithmetic of Power Monoids”, hosted throughout 2023 on the Art of Problem Solving (AoPS) website at
@@
As a non-trivial first step, Aggarwal et al. [1, Corollary 5.3] have shown by a probabilistic approach that, again in the case $H=\mathbb{N}$, the conjecture holds for almost all $l\in\mathbb{N}$, where “almost all” is understood in the sense of asymptotic density.